A company maucfactures batteries in batches of 14 and their is 3 percent reat of defects find the mean and standard deviation for the random variable x and the number of defects per bacth
Answer by stanbon(75887)

A company maucfactures batteries in batches of 14 and there is 3 percent risk of defects. find the mean and standard deviation for the random variable x and the number of defects per batch
--------
mean = np = 14*0.03 = 0.42
std = sqrt(npq) = sqrt(0.42*0.97) = 0.6383
--------------------------------------------
# of defects per batch = mean = 0.42
=======================================
Cheers,
Stan H.